### v3

# Hilbert-Kunz multiplicity of fiber products and idealization rings

Abstract.
The main purpose of this paper is to provide formulas for the Hilbert-Kunz multiplicity of fiber products and idealization rings. We calculate the Hilbert-Kunz multiplicity of a fiber product $R\times_{T}S$, with $R$, $S$ and $T$ being Noetherian local rings with common characteristic and residue field. Inspired by a work of Strazzanti and Armengou we calculate the Hilbert-Kunz multiplicity of an idealization ring $R\ltimes M$, where $M$ is a finite $R$-module. We explore certain examples and structural consequences and obtain some bounds for the Hilbert-Kunz multiplicity.
Key words and phrases:
Hilbert-Kunz, multiplicity, fiber product, idealization
2020 Mathematics Subject Classification:
13D40, 13H15, 13H05
Corresponding author: P. D. C. Martins
The first author was supported by grant 2019/21181-0, São Paulo Research Foundation (FAPESP). The second author was supported by grant 2022/12114-0, São Paulo Research Foundation (FAPESP)
1. Introduction
Throughout this paper, $(R,\mathfrak{m},k)$ will denote a Noetherian local ring of prime characteristic $p>0$ and dimension $d:=\dim R>0$. It’s worth noting that in this context, the residue field $k$ also shares the characteristic $p$. We will denote the length of an $R$-module by $\ell_{R}(-)$. The well-known Hilbert-Samuel multiplicity is a classical invariant of a local ring that quantifies singularity, where the minimum value of 1 corresponds to a smooth point. In the exploration of singularities, we naturally investigate equimultiple points or the equisingularity of deformations, concepts extensively studied in commutative algebra, algebraic geometry and singularity theory.
In 1983, Monsky [18] introduced a refined version of multiplicity specific to Noetherian local rings of prime characteristic, termed the Hilbert-Kunz multiplicity. It’s worth noting that this notion was also introduced by Kunz [16] in 1969, but Monsky provided a more explicit definition. While this invariant remains largely enigmatic, its study has revealed intriguing connections with the classical multiplicity, albeit often exhibiting more intricate behavior. Unlike the Hilbert-Samuel multiplicity, which is always an integer, the Hilbert-Kunz multiplicity may not be. However, this characteristic enables us to regard it as a more nuanced invariant. Similar to the classical multiplicity, it is capable of discerning regular rings. The study of Hilbert-Kunz multiplicity has engaged numerous researchers, as evidenced by works such as [18], [24], [8], and others.
1.1 Definition.
Let $I$ be an $\mathfrak{m}$-primary ideal of $R$ and let $M$ be a finite $R$-module. The Hilbert-Kunz multiplicity $e_{HK}(I,M)$ of $M$ with respect to $I$ is defined as follows:
(1.1) $e_{HK}(I,M):=\lim_{e\to\infty}\frac{\ell_{R}\left(M/I^{[q]}M\right)}{q^{d}},$
where $q=p^{e}$ and $I^{[q]}$ is the ideal generated by $x^{q}$, $x\in I$. If $M$ is a finite $R$-module $M/I^{[q]}M$ has finite lenght. For simplicity, we can put $e_{HK}(I):=e_{HK}(I,R)$ and $e_{HK}(R):=e_{HK}(\mathfrak{m})$, if it doesn’t cause confusion. Notice that the limit of the right-hand side of (1.1) always exists and is a real number; see, e.g., Monsky [18].
In [17], Kunz observed that the number $\ell_{e}(R):=\ell_{R}(R/\mathfrak{m}^{[q]})/q^{d}$ is a reasonable measure for the singularity of the local ring $R$ and gives a numerical measure of the ’singular-ness’ of $R$. In some sense, it is expected that more larger is the value of $e_{HK}(R)-1$, then more severe is the singularity of $R$.
If $R$ is regular, then $e_{HK}(I,R)=\ell_{R}(R/I)$ for every $\mathfrak{m}$-primary ideal $I$ of $R$ and $e_{HK}(R)=1$. Thus, it is natural to ask: ”If $e_{HK}(R)=1$, is $R$ then regular?”. There is a positive answer to the above question, given by Watanabe and Yoshida [23], when $R$ is unmixed, i. e., given an unmixed local ring $(R,\mathfrak{m})$ of characteristic $p>0$, if $e_{HK}(R)=1$, then $R$ is regular. A somewhat simpler proof can be found in [15]. Watanabe and Yoshida [24] also proved that it holds when $R$ is a two-dimensional Cohen-Macaulay local ring.
The main purpose of this paper is to provide new formulas for the Hilbert-Kunz multiplicity of fiber products and idealization rings and consequently produce criteria regarding the regularity of such rings. The study of fiber products and idealization rings has recently received special attention (see for instance [2, 7, 12, 13, 19, 20]). One reason to obtain formulas for the Hilbert-Kunz multiplicity of fiber product rings comes from the fact that the structure of the fiber product ring changes severely, depending on the rings considered (see Corollary 3.6 and Proposition 3.7).
Let us now briefly describe the contents of the paper. Section 2 establishes the main definitions and facts for the rest of the paper. Section 3 is devoted to study the Hilbert-Kunz multiplicity of fiber product rings. Some structural results and examples are also given. Motivated by a result of [21], whose obtain as a consequence a formula for the Hilbert-Kunz multiplicity for idealizations of the form $R\ltimes I$, where $I$ is an $\mathfrak{m}$-primary ideal of $R$, in the Section 4 we generalize and study the Hilbert-Kunz multiplicity of the idealization ring $R\ltimes M$, where $M$ is a finite $R$-module. Some structural results are provided. In the last section, we discuss two conjectures concerning lower bounds for the Hilbert-Kunz multiplicity. In particular, we provide new classes of $d$-dimensional Noetherian local rings of characteristic $p$ whose satisfy a conjecture posed by Watanabe and Yoshida [25], that asks if $e_{HK}(R)$ is bounded by $e_{HK}(R_{p,d})$, where $(R,\mathfrak{m},k)$ is a $d$-dimensional unmixed local ring of characteristic $p>2$ and dimension $d\geq 2$ and $R_{p,d}$ is the hypersurface $k[[X_{0},X_{1},\ldots,X_{d}]]/(X_{0}^{2}+\cdots+X_{d}^{2})$, with $k=\overline{\mathbb{F}_{p}}$.
2. Setup and Basic facts
We will start by recalling some fundamental properties of the Hilbert-Kunz multiplicity. After that, we will introduce the notion of fiber products and idealization rings to study their Hilbert-Kunz multiplicities. We observe that these two constructions work in general, when the ring characteristic is not necessarily positive.
Recall that if $R$ is a Noetherian local ring and $M$ is a finite $R$-module, then $\dim M$ is defined as $\dim M:=\dim(R/\operatorname{Ann}_{R}(M))$. Also, if $R$ is a local domain, then the rank of $M$ is $\operatorname{rk}_{R}M:=\dim_{Q}(Q\otimes_{R}M)$, where $Q$ is the total quotient ring of $R$.
2.1 Proposition (Fundamental properties (cf. [14, 16, 17, 18, 23])).
Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$. Let $I,J$ be $\mathfrak{m}$-primary ideals and let $M$ be a finite $R$-module. Then the following statements hold:
  • (1)
    If $I\subseteq J$, then $e_{HK}(I)\geq e_{HK}(J)$.
  • (2)
    $e_{HK}(I)\geq e_{HK}(R)\geq 1$.
  • (3)
    $e_{HK}(I,M)\geq 0$ and equality holds if and only if $\dim M<\dim R.$
  • (4)
    (Additivity in short exact sequences) If $0\longrightarrow L\longrightarrow M\longrightarrow N\longrightarrow 0$ is a short exact sequence of finite $R$-modules, then
    $e_{HK}(I,M)=e_{HK}(I,L)+e_{HK}(I,N).$
  • (5)
    (Associative Formula)
    $e_{HK}(I,M)=\sum_{\mathfrak{p}\in\operatorname{Assh}(R)}e_{HK}(I,R/\mathfrak{p})\cdot\ell_{R_{\mathfrak{p}}}(M_{\mathfrak{p}}).$
    where $\operatorname{Assh}(R)$ denotes the set of prime ideals $\mathfrak{p}$ of $R$ with $\dim R/\mathfrak{p}=\dim R$. In particular, if $R$ is a local domain and $M$ is a torsion-free $R$module of rank $\operatorname{rk}_{R}M$ , then $e_{HK}(I,M)=\operatorname{rk}_{R}M\cdot e_{HK}(R)$
  • (6)
    If $R$ is a local domain, then $e_{HK}(I,M)=\operatorname{rk}_{R}M\cdot e_{HK}(I,R)$.
  • (7)
    If $R$ is regular, then $e_{HK}(I)=\ell_{R}(R/I)$ and $e_{HK}(R)=1$.
  • (8)
    If $x\in I$ is $R$-regular, then $e_{HK}(I)\leq e_{HK}(I/xR)$.
  • (9)
    (Localization) $e_{HK}(R_{\mathfrak{p}})\leq e_{HK}(R)$ for any prime ideal $\mathfrak{p}$ such that $\dim R/\mathfrak{p}+\operatorname{height}\mathfrak{p}=\dim R$
  • (10)
    If $(A,\mathfrak{m})\to(B,\mathfrak{n})$ is a flat local ring homomorphism such that $B/\mathfrak{m}B$ is a field, then $e_{HK}(I)=e_{HK}(IB)$.
Fiber product rings
The fiber product structure is studied in various categories. In the case of Noetherian local rings, some fundamental properties can be found in [2].
2.2 Definition.
Let $(R,\mathfrak{m}_{R},k)$, $(S,\mathfrak{m}_{S},k)$ and $(T,\mathfrak{m}_{T},k)$ be Noetherian local rings, and let $R\stackrel{{\scriptstyle\pi_{R}}}{{\twoheadrightarrow}}T\stackrel{{\scriptstyle\pi_{S}}}{{\twoheadleftarrow}}S$ be surjective homomorphisms of rings. The fiber product
$R\times_{T}S=\{(r,s)\in R\times S\ \mid\ \pi_{R}(r)=\pi_{S}(s)\},$
is a Noetherian local ring with maximal ideal $\mathfrak{m}=\mathfrak{m}_{R}\times_{\mathfrak{m}_{T}}{\mathfrak{m}_{S}}$, residue field $k$, and it is a subring of the usual direct product $R\times S$ (see [2, Lemma 1.2]). Let $\eta_{R}:R\times_{T}S\twoheadrightarrow R$ and $\eta_{S}:R\times_{T}S\twoheadrightarrow S$ be the natural projections $(r,s)\mapsto r$ and $(r,s)\mapsto s$, respectively. Then $R\times_{T}S$ is represented as a pullback diagram:
2.3 Convention.
In general, when we use the symbols $R$, $S$ or $T$, we assume that $R$, $S$ and $T$ are non-zero Noetherian local rings with the maximal ideal $\mathfrak{m}_{R}$, $\mathfrak{m}_{S}$ and $\mathfrak{m}_{T}$, respectively, and with $k$ being the common residue field. Remark that each (finite) module over $R$, $S$ or $T$ acquires a canonical stucture of (finite) $R\times_{T}S$-module via Diagram (2.5).
We will state three already-known facts about fiber product rings to be used in Section 3.
2.4 Fact ([2]).
There is the exact short sequence of $R\times_{T}S$-modules
2.5 Fact ([2], Lemma 1.5).
The following (in)equalitie holds
$\dim(R\times_{T}S)=\max\{\dim R,\dim S\}\geq\min\{\dim R,\dim S\}\geq\dim T$
2.6 Fact.
Let $k$ be a field and let $I$ and $J$ be ideals of $k[[X_{1},\dots,X_{n}]]$ and $k[[Y_{1},\dots,Y_{m}]]$, respectively. Then
$\frac{k[[X_{1},\dots,X_{n}]]}{I}\times_{k}\frac{k[[Y_{1},\dots,Y_{m}]]}{J}\cong\frac{k[[X_{1},\dots,X_{n},Y_{1},\dots,Y_{m}]]}{(I+J+(X_{i}Y_{j}:1\leq i\leq n,1\leq j\leq m))}.$
Idealization of a module
Next construction was introduced by Nagata and its nice properties were studied by Anderson and Winders [3].
2.7 Definition.
Let $R$ be a commutative ring and let $M$ be a $R$-module. Then $R\ltimes M:=R\times M$ with coordinate wise addition and multiplication given by $(r_{1},m_{1})(r_{2},m_{2}):=(r_{1}r_{2},r_{1}m_{2}+r_{2}m_{1})$ is a commutative ring with unity (1,0) called the idealization of $M$.
2.8 Remark.
Taking a finite $R$-module $M$ over a $d$-dimensional Noetherian local ring of prime chatecteristic $p>0$ with maximal ideal $\mathfrak{m}$ and residue field $k$, it is well known that the idealization ring $R\ltimes M$ is a $d$-dimensional Noetherian local ring of positive chatecteristic $p$ with maximal ideal $\mathfrak{m}\ltimes M$ and residue field k. Also, we can consider $R\hookrightarrow R\ltimes M$ by sending $r\mapsto(r,0)$. Then $R\ltimes M$ is a local $R$-algebra with the same residue field of R. It is discussed in [3].
We will state two already proved theorems about idealization rings to be used in Section 4.
2.9 Theorem ([3], Theorem 3.1).
Let $R$ be a commutative ring, $I$ an ideal of $R$, M an $R$-module and $N$ a submodule of $M$. Then $I\ltimes N$ is an ideal of $R\ltimes M$ if and only if $IM\subseteq N$. When $I\ltimes N$ is an ideal, $M/N$ is an $R/I$-module and $(R\ltimes M)/(I\ltimes M)\cong(R/I)\ltimes(M/N)$. In particular $(R\ltimes M)/(0\ltimes N)\cong R\ltimes(M/N)$ and therefore $(R\ltimes M)/(0\ltimes M)\cong R$. So the ideals of $R\ltimes M$ containing $0\ltimes M$ are of the form $J\ltimes M$ for some ideal $J$ of $R$.
2.10 Theorem ([3], Theorem 3.6).
Let $R$ be a commutative ring and $M$ an $R$-module. Let $I$ be an ideal of $R$ and $N$ a submodule of $M$. Then $I\ltimes N$ is primary if and only if either
  • a)
    $N=M$ and $I$ is a primary ideal of $R$ or
  • b)
    $N\subsetneq M$, $IM\subseteq N$, and $I$ and $N$ are $\mathfrak{p}$-primary where $\mathfrak{p}=\sqrt{I}$
In either case, $I\ltimes N$ is $\sqrt{I}\ltimes M$-primary.
3. Hilbert-Kunz Multiplicity of Fiber Product Rings
In this section, the rings $(R,\mathfrak{m}_{R},k)$, $(S,\mathfrak{m}_{S},k)$ and $(T,\mathfrak{m}_{T},k)$ will be considered Noetherian local rings as in the Definition 2.2. Also, assume that $R$ and $S$ have positive characteristic $p$ and $q$, respectively. In this case, the fiber product ring $R\times_{T}S$ is also a Noetherian local ring of positive characteristic $\max\{p,q\}$. During this section, fiber product rings $R\times_{T}S$ will be considered under these conditions. To not be confused, note that in this section the rings $R$, $S$, and $T$ will also be considered as $R\times_{T}S$-modules.
3.1 Lemma.
Let $R\times_{T}S$ be a Noetherian local ring of characteristic $p>0$ and let $I$ be an $\mathfrak{m}$-primary ideal of $R\times_{T}S$. Then the following equality holds
$e_{HK}(I)=\left\{\begin{array}[]{llll}e_{HK}(I,R)+e_{HK}(I,S)-e_{HK}(I,T);&\dim R=\dim S=\dim T\\ e_{HK}(I,R)+e_{HK}(I,S);&\dim R=\dim S>\dim T\\ e_{HK}(I,R);&\dim R>\dim S>\dim T.\end{array}\right.$
Proof.
Using the short exact sequence of $R\times_{T}S$-modules given in Fact 2.4 and the additivity of Hilbert-Kunz multiplicity in exact short sequences, we get
$e_{HK}(I)=e_{HK}(I,R\oplus S)-e_{HK}(I,T).$
Applying the same argument to the exact sequence
we obtain $e(I,R\oplus S)=e(I,R)+e(I,S)$ and then
$e_{HK}(I)=e_{HK}(I,R)+e_{HK}(I,S)-e_{HK}(I,T)$
Note that the dimensions of $R$, $S$ and $T$ as $R\times_{T}S$-modules coincide with their respective Krull dimensions. Finally, by using the formula for $\dim(R\times_{T}S)$ given in Fact 2.5 and item (3) of Proposition 2.1 we get the desired formulas. ∎
Next proposition provides a formula for the Hilbert-Kunz multiplicity of a fiber product $R\times_{T}S$ in terms of the Hilbert-Kunz multiplicities of $R$, $S$ and $T$, when $R$, $S$ and $T$ have the same characteristic.
3.2 Proposition.
Let $R$, $S$ and $T$ be Noetherian local rings of characteristic $p>0$ as in (2.5). Then the following equality holds
$e_{HK}(R\times_{T}S)=\left\{\begin{array}[]{llll}e_{HK}(R)+e_{HK}(S)-e_{HK}(T);&\dim R=\dim S=\dim T\\ e_{HK}(R)+e_{HK}(S);&\dim R=\dim S>\dim T\\ e_{HK}(R);&\dim R>\dim S>\dim T.\end{array}\right.$
Proof.
As an immediate consequence of Lemma 3.1, we obtain the formulas
(3.4) $e_{HK}(R\times_{T}S)=\left\{\begin{array}[]{llll}e_{HK}(\mathfrak{m},R)+e_{HK}(\mathfrak{m},S)-e_{HK}(\mathfrak{m},T);&\dim R=\dim S=\dim T\\ e_{HK}(\mathfrak{m},R)+e_{HK}(\mathfrak{m},S);&\dim R=\dim S>\dim T\\ e_{HK}(\mathfrak{m},R);&\dim R>\dim S>\dim T.\end{array}\right.$
Now, we will discuss why we can replace $e_{HK}(\mathfrak{m},R)$, $e_{HK}(\mathfrak{m},S)$ and $e_{HK}(\mathfrak{m},T)$ in (3.4) by $e_{HK}(R)$, $e_{HK}(S)$ and $e_{HK}(T)$, respectively. Note that $R\times_{T}S$ also has characteristic $p$. In general, if $(A,\mathfrak{m}_{A},k)$ and $(B,\mathfrak{m}_{B},k)$ are $d$-dimensional Noetherian local rings of characteristic $p>0$ and $f:A\to B$ is a surjective ring homomophism, then $f(\mathfrak{m}_{A})=\mathfrak{m}_{B}$ and so it is easy to see that $\mathfrak{m}_{A}^{[q]}B=\mathfrak{m}_{B}^{[q]}B$, with $q=p^{e}$, for each positive integer $e$. Also, as $\ell_{A}(N)=\ell_{B}(N)$, for all $B$-module $N$, we obtain:
$e_{HK}(\mathfrak{m}_{A},B)$ $=\lim_{e\to\infty}\frac{\ell_{A}\left(B/\mathfrak{m}^{[q]}_{A}B\right)}{q^{d}}$
$=\lim_{e\to\infty}\frac{\ell_{B}\left(B/\mathfrak{m}_{B}^{[q]}\right)}{q^{d}}$
$=e_{HK}(B).$
Therefore, by utilizing the formula for $\dim R\times_{T}S$ provided in Fact 2.5, the surjective ring homomorphisms of the diagram (2.5), and equality (3.4), we obtain the desired formulas, by observing that we need only to consider the components of the fiber product $R\times_{T}S$ whose have dimension equal to $\dim R\times_{T}S$. ∎
3.1 Importance of the formula.
Watanabe and Yoshida asked whether it is possible to determine the infimum of Hilbert-Kunz multiplicities over all unmixed Noetherian local rings of characteristic $p>0$ and dimension $d$ that are not regular. Due to this and the Watanabe-Yoshida conjecture (see Conjecture 5.3), some authors (see, e.g., [25, 6, 4, 1]) have determined lower and upper bounds for some unmixed local rings. However, it is also natural to ask whether it is possible to find lower and upper bounds for the Hilbert-Kunz multiplicity of Noetherian local rings that are not necessarily unmixed. In this sense, by using existing results, Proposition 3.2 helps us to find bounds for the fiber product ring (which is not always unmixed) by knowing the bounds of the components of the fiber product ring. Below, we provide one example among many to illustrate this fact:
3.3 Example.
Let $R$, $S$, and $T$ be Noetherian local rings of characteristic $p>0$ as in (2.5) with $d:=\dim(R)\geq 2$ and $\dim R\geq\dim S>\dim T$. Suppose that $R$ is a formally unmixed local ring that is not regular. Then
$e_{HK}(R\times_{T}S)\geq 1+\frac{1}{d!(d(d+1)+1)^{d}}.$
The inequality follows easily from Proposition 3.2 and [1, Theorem 4.12].
Amalgamated duplication
Let $R$ be a Noetherian local ring. If we assume $\pi_{R}=\pi_{S}:R\to R/I$ (Diagram (2.5)) are the canonical maps and $I$ is a proper $R$-ideal, there is an isomorphism $R\times_{R/I}R\cong R\bowtie I$, where $R\bowtie I$ is the amalgamated duplication of $R$ along $I$, introduced by D’Anna [9]. Remark that, if $R$ is local with maximal ideal $\mathfrak{m}$, then $R\bowtie I$ is a local ring with maximal ideal denoted by $\mathfrak{m}\bowtie I.$ Some structural properties of $R\bowtie I$ are summarized in [2, Theorem 1.8 and 1.9 - 1.9.4].
We can use Proposition 3.2 to obtain a formula for the Hilbert-Kunz Multiplicity of the amalgamated duplication along an ideal.
3.4 Corollary.
Let $R$ be a Noetherian local ring of characteristic $p>0$ and let $I$ be an ideal of $R$. If $\dim R/I<\dim R$, then
$e_{HK}(R\bowtie I)=2e_{HK}(R).$
Proof.
It follows directly by Proposition 3.2. ∎
3.5 Remark.
The formula given in Corollary 3.4 generalizes a formula obtained in [21, Theorem 2.3 and Remark 2.4] for the Hilbert-Kunz multiplicity of an amalgameted duplication of $R$ along an $\mathfrak{m}$-primary ideal $I$. In this case $\dim R/I=0$ and then $e_{HK}(R\bowtie I)=2e_{HK}(R)$.
The study of structural properties of fiber product rings has been addressed by several authors, as [2] and [20]. Next corollary provides some structural properties about fiber product rings.
3.6 Corollary.
Let $R$, $S$ and $T$ be Noetherian local rings of characteristic $p>0$ as in (2.5) and let $I$ be an ideal of $R$. Then
  • (1)
    If $\dim R=\dim S=\dim T$ and $e_{HK}(R)+e_{HK}(S)-e_{HK}(T)>1$, then $R\times_{T}S$ is not a regular local ring.
  • (2)
    If $\dim R=\dim S>\dim T$, then $R\times_{T}S$ is not a regular local ring. In particular, if $\dim R>\dim R/I$, then $R\bowtie I$ is not a regular local ring.
  • (3)
    If $\dim R>\dim S>\dim T$ and $e_{HK}(R)>1$, then $R\times_{T}S$ is not a regular local ring.
Proof.
It is well known that regular local rings have Hilbert-Kunz multiplicity equal to one. In all the above-mentioned cases, using Proposition 3.2, we obtain $e_{HK}(R\times_{T}S)>1$ and then $R\times_{T}S$ can not be regular. ∎
As we showed above, in general, fiber product rings are not regular. The next result provides an equivalent condition for some fiber product rings to be regular.
3.7 Proposition.
Let $R$, $S$ and $T$ be Noetherian local rings of characteristic $p>0$ as in (2.5). If one of the following conditions holds:
  • (1)
    $R$, $S$ and $T$ are two-dimensional Cohen-Macaulay local rings.
  • (2)
    $R\times_{T}S$ is unmixed.
Then $R\times_{T}S$ is regular if and only if $e_{HK}(R)+e_{HK}(S)-e_{HK}(T)=1.$
Proof.
  • (1)
    By Proposition 3.2, we have $e_{HK}(R\times_{T}S)=e_{HK}(R)+e_{HK}(S)-e_{HK}(T)$. Also, we observe that as $R$, $S$ and $T$ are two-dimensional Cohen-Macaulay local rings, then the same is true for $R\times_{T}S$, by [2, Proposition 1.7]. For two-dimensional Cohen-Macaulay local rings, regularity is equivalent to unitary Hilbert-Kunz multiplicity, by [24, Theorem 2.5]). Then $R\times_{T}S$ is regular if and only if
    $e_{HK}(R\times_{T}S)=e_{HK}(R)+e_{HK}(S)-e_{HK}(T)=1.$
  • (2)
    The proof follows by Proposition 3.2 and [23, Theorem 1.5].
3.8 Corollary.
Let $R$, $S$ and $T$ be two-dimensional regular local rings of characteristic $p>0$ as in (2.5). Then $R\times_{T}S$ is regular.
To end this section, we will use the formulas given in Proposition 3.2 to calculate the Hilbert-Kunz multiplicity of some fiber product rings.
3.9 Example.
Let $k$ be a field of characteristic $p>0$. If $R=k[[X_{1},\dots,X_{n}]]$ and $S=k[[Y_{1},\dots,Y_{m}]]$, then
$R\times_{k}S\cong\frac{k[[\underline{X},\underline{Y}]]}{(X_{i}Y_{j}:1\leq i\leq n,1\leq j\leq m)},$
by Fact 2.6. Therefore, if $n>m$, then $\dim R>\dim S>\dim k$ and we have
(3.5) $e_{HK}\left(\frac{k[[\underline{X},\underline{Y}]]}{(X_{i}Y_{j}:1\leq i\leq n,1\leq j\leq m)}\right)=e_{HK}(K[[\underline{X}]])=1,$
by Proposition 3.2. Note that (3.5) provides examples of non-regular local rings with Hilbert-Kunz multiplicity $1$. Also, if $n=m$, then $\dim R=\dim S>\dim k$ and we have
$e_{HK}\left(\frac{k[[\underline{X},\underline{Y}]]}{(X_{i}Y_{j}:1\leq i,j\leq n)}\right)=e_{HK}(K[[\underline{X}]])+e_{HK}(K[[\underline{Y}]])=2,$
by Proposition 3.2.
3.10 Example.
Let $k$ be a field of characteristic $p\geq 3$. Consider the hypersurfaces and $R=k[[X,Y,Z]]/(XY+Z^{n+1})$ and $S=k[[W,U,V]]/(W^{2}+UV^{2}+U^{n-1})$, where $n\in\mathbb{N}$. Then
$R\times_{k}S\cong\frac{k[[X,Y,Z,W,U,V]]}{(XW,XU,XV,YW,YU,YV,ZW,ZU,ZV,XY+Z^{n+1},W^{2}+UV+U^{n-1})},$
by Fact 2.6. In this case, $\dim R=\dim S>\dim k$ and by using the formulas given in [14, Example 3.18], we have
$e_{HK}(R\times_{k}S)$ $=e_{HK}(R)+e_{HK}(S)=\left(2-\frac{1}{n+1}\right)+\left(2-\frac{1}{4(n-2)}\right)$
$=4-\frac{5n-7}{4(n+1)(n-2)},$
for all positive integer $n\geq 4$, by Proposition 3.2.
3.11 Example.
Let $k$ be a field of characteristic $p>0$. Let $R=k[[X]]$, $S=k[[Y]]$ and $T=K[[Z]]/(Z^{n})$, where $n\in\mathbb{N}$ and $n\geq 2$. Then
$R\times_{T}S\cong\frac{k[[X,Y]]}{(X^{n}Y-Y^{2})},$
by [5, Corollary 4.2.7], taking $A=k[[X,Y]]/(X^{n}Y-Y^{2})$, $\mathfrak{p}=(X^{n}-Y)$ and $\mathfrak{q}=(Y)$. In this case, $\dim R=\dim S>\dim T$ and we have
$e_{HK}\left(\frac{k[[X,Y]]}{(X^{n}Y-Y^{2})}\right)=e_{HK}(k[[X]])+e_{HK}(k[[Y]])=2,$
for all positive integer $n\geq 2$, by Proposition 3.2.
4. Hilbert-Kunz multiplicity of idealization rings
We will start by proving the following lemma.
4.1 Lemma.
Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be an $R$-module. If $I$ is an ideal of $R$, $N$ is a submodule of $M$ and $J=I\ltimes N$ is an ideal of $R\ltimes M$, or equivalently $IM\subseteq N$, then
$J^{[p^{e}]}=I^{[p^{e}]}(R\ltimes M),$
for all positive integer $e$.
Proof.
First of all, we will fix $q:=p^{e}$. Note that, for $(r,m)\in R\ltimes M$ we have that $(r,m)^{n}=(r^{n},nr^{n-1}m)$, for all positive integer $n$. So
$(r,m)^{q}=(r^{q},qr^{q-1}m)=(r^{q},p^{e}r^{q-1}m)=(r^{q},0).$
Then the generators of the ideal $J^{[q]}$ are of the form $(r^{q},0)$, where $r\in I$, and its elements have the form $\sum_{i=1}^{n}(s_{i},m_{i})(r_{i}^{q},0)$ where $r_{i}\in I$, $s_{i}\in R$ and $m_{i}\in M$, for $i=1,\dots,n$. Observe that
$\sum_{i=1}^{n}(s_{i},m_{i})(r_{i}^{q},0)=\sum_{i=1}^{n}(r_{i}^{q},0)(s_{i},m_{i})\in I^{[q]}(R\ltimes M).$
Then $J^{[q]}\subseteq I^{[q]}(R\ltimes M)$. On other hand, the generators of the $R$-module $I^{[q]}(R\ltimes M)$ are of the form $r(s,m)$, where $r\in I^{[q]}$ and $(s,m)\in R\ltimes M$. Writing $r=\sum_{i=1}^{n}y_{i}x_{i}^{q}$, where $y_{i}\in R$ and $x_{i}\in I$, for $i=1,\dots,n$, we get
$r(s,m)=\sum_{i=1}^{n}y_{i}x_{i}^{q}(s,m)=\sum_{i=1}^{n}(y_{i}s,y_{i}m)(x_{i}^{q},0)\in J^{[q]}.$
Then $I^{[q]}(R\ltimes M)\subseteq J^{[q]}$ and the equality holds. ∎
The next theorem provides as a consequence a formula for the Hilbert-Kunz multiplicity of an idealization ring $R\ltimes M$, where $M$ is a finite $R$-module.
4.2 Theorem.
Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finite $R$-module. If one of the following conditions holds:
  • (1)
    $I$ is an $\mathfrak{m}$-primary ideal of $R$ and $J=I\ltimes M$.
  • (2)
    $I$ is an $\mathfrak{m}$-primary ideal of $R$, $N\subsetneq M$ is a submodule of $M$ which is $\mathfrak{m}$-primary and such that $J=I\ltimes N$ is an ideal of $R\ltimes M$ , or equivalently $IM\subseteq N$.
Then following equalities holds
$e_{HK}(J,R\ltimes M)=e_{HK}(I,R)+e_{HK}(I,M)=\lim_{e\to\infty}\frac{\ell_{R}(R\ltimes M/I^{[q]}\ltimes I^{[q]}M)}{q^{d}}.$
In particular, if $\dim M<\dim R$, then $e_{HK}(J,R\ltimes M)=e_{HK}(I,R)$.
Proof.
Note that in both cases the ideal $J$ is $\mathfrak{m}\ltimes M$-primary, by Theorem, 2.10. Fix $q:=p^{e}$. Using the canonical short exact sequence of $R$-modules
we get
$e_{HK}(I,R\ltimes M)=e_{HK}(I,R)+e_{HK}(I,M).$
Observe that $\ell_{R}(P)=\ell_{R\ltimes M}(P)$, for each $R\ltimes M$-module $P$. Using Lemma 4.1 we have $J^{[q]}=I^{[q]}(R\ltimes M)$ and then
$e_{HK}(J,R\ltimes M)$ $=\lim_{e\to\infty}\frac{\ell_{R\ltimes M}(R\ltimes M/J^{[q]})}{q^{d}}$
$=\lim_{e\to\infty}\frac{\ell_{R}(R\ltimes M/I^{[q]}(R\ltimes M))}{q^{d}}$
$=e_{HK}(I,R\ltimes M)$
$=e_{HK}(I,R)+e_{HK}(I,M).$
Also, by Theorem 2.9, we obtain the isomorphism of $R$-algebras
$\frac{R\ltimes M}{I^{[q]}\ltimes I^{[q]}M}\cong\frac{R}{I^{[q]}}\ltimes\frac{M}{I^{[q]}M}$
and an exact sequence of $R$-modules
[…]
Using length additivity in short exact sequences we obtain
$\ell_{R}\left(\frac{R\ltimes M}{I^{[q]}\ltimes I^{[q]}M}\right)=\ell_{R}\left(\frac{R}{I^{[q]}}\right)+\ell_{R}\left(\frac{M}{I^{[q]}M}\right).$
Now, taking limits we get the equality:
$\lim_{e\to\infty}\frac{\ell_{R}(R\ltimes M/I^{[q]}\ltimes I^{[q]}M)}{q^{d}}$ $=\lim_{e\to\infty}\frac{\ell_{R}(R/I^{[q]})}{q^{d}}+\lim_{e\to\infty}\frac{\ell_{R}(M/I^{[q]}M)}{q^{d}}$
$=e_{HK}(I,R)+e_{HK}(I,M).$
4.3 Corollary.
Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finite $R$-module. Then
  • (1)
    $e_{HK}(R\ltimes M)=e_{HK}(R)+e_{HK}(\mathfrak{m},M)$. In particular, if $\dim M<\dim R$, then $e_{HK}(R\ltimes M)=e_{HK}(R)$.
  • (2)
    If $I$ is an ideal of $R$ and $\dim R/I<\dim R$, then $e_{HK}(R\ltimes I)=2e_{HK}(R)$.
Proof.
  • (1)
    It is immediate from Theorem 2.9, by taking $J=\mathfrak{m}\ltimes R$.
  • (2)
    Using the canonical short exact sequence of $R$-modules
    we get $e_{HK}(R)=e_{HK}(\mathfrak{m},I)+e_{HK}(\mathfrak{m},R/I)$. Since $\dim R/I<\dim R$, then $e_{HK}(\mathfrak{m},R/I)=0$ and $e_{HK}(R)=e_{HK}(\mathfrak{m},I)$. Utilizing item (1) we obtain the desired equality.
4.4 Remark.
The formulas given in Corollary 4.3 generalize a formula obtained in [21, Theorem 2.3 and Remark 2.4] for the Hilbert-Kunz multiplicity of an idealization ring $R\ltimes I$, where $I$ is an $\mathfrak{m}$-primary ideal of $R$. In this case $\dim R/I=0$ and then $e_{HK}(R\ltimes I)=2e_{HK}(R)$.
4.5 Corollary.
Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finite $R$-module. Then
  • (1)
    If $\dim M=\dim R$ or $e_{HK}(R)>1$, then $R\ltimes M$ is not a regular local ring.
  • (2)
    If $R$ is unmixed and there exists a finite $R$-module $M$ with $e_{HK}(R\ltimes M)=1$, then $\dim M<\dim R$ and $R$ is a regular local ring.
Proof.
  • (1)
    By Corollary 4.3, we have $e_{HK}(R\ltimes M)=e_{HK}(R)+e_{HK}(\mathfrak{m},M)>1,$ as we suppose $\dim M=\dim R$ or $e_{HK}(R)>1$. Then $R\ltimes M$ can not be regular.
  • (2)
    Using Corollary 4.3, we get $e_{HK}(R)=1$ and $e_{HK}(\mathfrak{m},M)=0$. As $R$ is unmixed it must be regular, by [23, Theorem 1.5]. Also, we have $\dim M<\dim R$, by item (3) of Proposition 2.1.
The next corollary concerns providing formulas for the Hilbert-Kunz multiplicity $e_{HK}(J,R\ltimes M)$, where $J$ is an ideal of $R\ltimes M$ as in Theorem 2.9, by using structural and homological properties of $M$.
4.6 Corollary.
Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finite $R$-module. Let $I$ be an $\mathfrak{m}$-primary ideal of $R$ and let $J$ be an ideal as in Theorem 2.9. Then
  • (1)
    $e_{HK}(J,R\ltimes M)\leq(1+\mu(M))e_{HK}(I,R)$.
  • (2)
    If $\operatorname{pd}_{R}M=n<\infty$, then
    $e_{HK}(J,R\ltimes M)=\left(\sum_{i=0}^{n}(-1)^{i}\beta_{i}^{R}(M)+1\right)e_{HK}(I,R),$
    where $\beta_{i}^{R}(M)$ denotes the i-th Betti number of $M$. In particular, if $M$ is a free $R$-module of rank $r$, then $e_{HK}(J,R\ltimes M)=(r+1)e_{HK}(I,R)$.
  • (3)
    If $R$ is a local domain, then
    $e_{HK}(J,R\ltimes M)=(\operatorname{rk}_{R}(M)+1)e_{HK}(I,R),$
    Also, if in addition we assume that $M$ is a torsion-free $R$-module of rank $r$, then $e_{HK}(J,R\ltimes M)=e_{HK}(I,M)+re_{HK}(R)$.
Proof.
  • (1)
    Considering a short exact sequence
    we get
    $e_{HK}\left(I,R^{\mu(M)}\right)=\mu(M)e_{HK}(I,R)=e_{HK}(I,K)+e_{HK}(I,M).$
    Then $\mu(M)e_{HK}(I,R)\geq e_{HK}(I,M)$. From this inequality and by Theorem 4.2, we obtain the result.
  • (2)
    Taking a minimal free-resolution of $M$
    and using the additivity of Hilbert-Kunz multiplicity on short exact sequences of finite $R$-modules, we obtain
    $e_{HK}(I,M)=\sum_{i=0}^{n}(-1)^{i}\beta_{i}^{R}(M)e_{HK}(I,R).$
    Utilizing Theorem 2.9 we obtain the desired equality.
  • (3)
    The proof follows by Theorem 2.9 and itens (5),(6) of Proposition 2.1.
4.7 Corollary.
Let $R$ be a Noetherian local ring of dimension $d>0$ and characteristic $p>0$. Then
$e_{HK}\left(\frac{R[X_{1},\dots,X_{r}]}{(X_{1},\dots,X_{r})^{2}}\right)=(r+1)e_{HK}(R).$
Proof.
If $M=R^{r}$, then $R\ltimes M$ is naturally isomorphic to $R[X_{1},\dots,X_{r}]/(X_{1},\dots,X_{r})^{2}$, by [3, Proposition 2.2]. Therefore, the equality follows by the above corollary. ∎
5. Some bounds for the Hilbert-Kunz multiplicity
There are some fundamental conjectures about Hilbert-Kunz multiplicity to be answered. The following conjecture of Watanabe and Yoshida can be found in [23].
5.1 Conjecture.
Let $R$ be a Cohen-Macaulay local ring of characteristic $p>0$. Then for any $\mathfrak{m}$-primary ideal $I$, we have
  • (1)
    $e_{HK}(I,R)\geq\ell_{R}(R/I)$.
  • (2)
    If $\operatorname{pd}_{R}(R/I)<\infty$, then $e_{HK}(I,R)=\ell_{R}(R/I)$.
It is well known that the above conjecture holds if $R$ is a complete intersection (see Dutta [10, Theorem 1.9]).
The next proposition extends the result already proved for complete intersections to an idealization ring $R\ltimes M$ where $M$ is a finite $R$-module, by using Theorem 4.2. It consists of obtaining an inequality as in the item (1) of Conjecture 5.1 proposed by Watanabe and Yoshida.
5.2 Proposition.
Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finite $R$-module. If $R$ is a complete intersection, $I$ is an $\mathfrak{m}$-primary ideal of $R$ and $J=I\ltimes M$, then
$e_{HK}(J,R\ltimes M)\geq\ell_{R\ltimes M}\left(\frac{R\ltimes M}{J}\right).$
Proof.
Since $R$ is a complete intersection, then it is true that $e_{HK}(I,R)\geq\ell_{R}(R/I)$, by [10, Theorem 1.9]. Using Theorem 4.2, we obtain
$e_{HK}(J,R\ltimes M)\geq e_{HK}(I,R)\geq\ell_{R}(R/I)$
and by Theorem 2.9 we have a canonical isomorphism $R\ltimes M/J\cong R/I$. Therefore, we have
$e_{HK}(J,R\ltimes M)\geq\ell_{R}\left(\frac{R}{I}\right)=\ell_{R\ltimes M}\left(\frac{R\ltimes M}{J}\right).$
New estimates on Watanabe-Yoshida Conjecture
As we observe in Section 1, it is well known that study of Hilbert-Kunz multiplicity is an important tool to investigate the regularity and singularity of unmixed local rings. In this sense, it is natural to ask:
Question: Is there a constant $\epsilon(d)$, depending only on the dimension $d$ of $R$, such that if $R$ is an unmixed local ring which is not regular, then $e_{\text{HK}}(R)-1\geq\epsilon(d)$? Aberbach and Enescu provided a positive answer to this question ([1]). However, if we remove the unmixed condition, does the answer remain valid?
The question above was made even more explicit when Watanabe and Yoshida posed the following conjecture.
5.3 Conjecture (Watanabe-Yoshida [25]).
Let $d\geq 2$ be an integer and $p>2$ be a prime number. Set
$R_{p,d}:=\frac{k[[X_{0},X_{1},\ldots,X_{d}]]}{(X_{0}^{2}+\cdots+X_{d}^{2})}.$
Let $(R,\mathfrak{m},k)$ be a $d$-dimensional unmixed local ring of characteristic $p$ and let $k=\overline{\mathbb{F}_{p}}$. Then the following statements hold:
  1. (1)
    If $R$ is not regular, then $e_{\text{HK}}(R)\geq e_{\text{HK}}(R_{p,d})\geq 1+m_{d}$.
  2. (2)
    If $e_{\text{HK}}(R)=e_{\text{HK}}(R_{p,d})$, then the $\mathfrak{m}$-adic completion of $R$ is isomorphic to $R_{p,d}$ as local rings.
Here, the numbers $m_{d}$ are defined by the relation
$\sec x+\tan x=1+\sum_{d=1}^{\infty}m_{d}x^{d}\quad(-\pi/2<x<\pi/2).$
Recently, Cox and Aberbach [8] proved that the Watanabe-Yoshida conjecture holds up to dimension $7$. The first inequality of part (1) of the conjecture is known in full generality for complete intersections by the work of Enescu and Shimomoto [11]. Also, recently, Trivedi [22] proved the second inequality of part (1) of the conjecture, for $d\geq 4$ and $p>d-3$.
In our work, we calculated their Hilbert-Kunz multiplicity and gave conditions for fiber products and idealization rings not to be regular. The next results show new classes of $d$-dimensional Noetherian local rings of characteristic $p>0$ whose satisfy the part (1) of Conjecture of Watanabe and Yoshida by using the formulas presented in the sections 3 and 4.
5.4 Proposition.
Let $d\geq 2$ be an integer and let $p>2$ be a prime number. Let $R$, $S$, and $T$ be Noetherian local rings of characteristic $p$ as in (2.5) and assume $R$ is a $d$-dimensional complete intersection, not regular. If $\dim R\geq\dim S>\dim T$, then
$e_{HK}(R\times_{T}S)\geq e_{HK}(R_{d,p}).$
If in addition $d\geq 4$ and $p>d-3$, then $e_{HK}(R\times_{T}S)\geq e_{HK}(R_{d,p})\geq 1+m_{d}$ and item (1) of Conjecture 5.3 holds for the fiber product $R\times_{T}S$.
Proof.
The proof follows by Proposition 3.2, [11, Theorem 4.6] and [22, Theorem (A)]. ∎
5.5 Corollary.
Let $d\geq 2$ be an integer, $p>2$ be a prime number, $(R,\mathfrak{m},k)$ be a $d$-dimensional complete intersection of characteristic $p$, not regular, and let $I$ be an ideal of $R$. If $\dim R/I<\dim R$, then
$e_{HK}(R\bowtie I)\geq e_{HK}(R_{d,p}).$
If in addition $d\geq 4$ and $p>d-3$, then $e_{HK}(R\bowtie I)\geq e_{HK}(R_{d,p})\geq 1+m_{d}$ and item (1) of Conjecture 5.3 holds for the amalgameted duplication $R\bowtie I$.
5.6 Proposition.
Let $d\geq 2$ be an integer, $p>2$ be a prime number and let $(R,\mathfrak{m},k)$ be a $d$-dimensional complete intersection of characteristic $p$, not regular. If $M$ is a finite $R$-module, then
$e_{HK}(R\ltimes M)\geq e_{HK}(R_{d,p}).$
If in addition $d\geq 4$ and $p>d-3$, then $e_{HK}(R\ltimes M)\geq e_{HK}(R_{d,p})\geq 1+m_{d}$ and item (1) of Conjecture 5.3 holds for the idealization ring $R\ltimes M$.
Proof.
The proof follows by item (1) of Corollary 4.3, [11, Theorem 4.6] and [22, Theorem (A)]
5.7 Remark.
Propositions 5.2 and 5.4, 5.6 effectively expand the class of rings and ideals that satisfy the desired inequalities of conjectures 5.1 and 5.3, respectively, because an idealization ring $R\ltimes M$ and a fiber product ring $R\times_{T}S$ may not even be Cohen-Macaulay when $R$ is a complete intersection local ring. Both situations are exemplified below.
5.8 Example ([3], Example 4.15).
Let $(R,\mathfrak{m})$ be an $d$-dimensional regular local ring with $x_{1},\dots,x_{d}$ a minimal basis for $\mathfrak{m}$. Let $M_{i}=R/(x_{1},\dots,x_{i})$, $0\leq i\leq d$ (for $i=0$, $M_{i}=R$). Then $\dim R\ltimes M_{i}=d$ and $\operatorname{depth}R\ltimes M_{i}=d-i$. So $R\ltimes M_{i}$ is not Cohen-Macaulay, for $0<i\leq d$.
5.9 Example.
Let $(R,\mathfrak{m}_{R},k)$ and $(S,\mathfrak{m}_{S},k)$ be Noetherian local rings and suppose $\dim R,\dim S\geq 2$. Then $R\times_{k}S$ is not Cohen-Macaulay, by Fact 2.5 and [7, Remark 3.2].
Competing interests: The authors declare none.